# GnnMerge: Merging of Gnn Models Without Accessing Training Data

### v1

@@ Appendix C Additional Results and Analysis
Tables 8910 and 11 contain the full results for the Different Datasets experiments, with varying number of models being merged.
@@ C.2 Generalization to GNN Architectures
Table 12 contains the results for In-domain Dataset experiments on GraphSage and NodeFormer. These further establish the robustness of GnnMerge to merge diverse GNN architectures.
C.3 Data Efficiency
C.3.1 Target Node Sampling
@@
Figure 6a and 6b depict the variation in test accuracies as the percentage of nodes utilized for alignment is varied on the arxiv and reddit datasets.
@@
Figure 6:
C.3.2 1-Hop Neighbour Condensation
In an Mpnn architecture, each layer computes node embeddings by aggregating messages from a node’s 1-hop neighbors. As a result, at any given layer, a target node’s embedding depends exclusively on its immediate neighbors, if the input is fixed. This property enables the graph to be reduced to only the target nodes and their 1-hop neighbors, significantly decreasing its size. When combined with node sampling, this leads to a substantial reduction in both memory requirements and merging time. Figures 7a and 7b illustrate the variation in memory consumption across different node sampling levels for the Arxiv and Reddit datasets, respectively. The overall sampling procedure leads to 3 benefits:
@@
(a) Graph Size(Edges) of Arxiv as the sampling ratio is varied
(b) Graph Size(Edges) of Reddit as the sampling ratio is varied
Figure 7:
@@ Appendix D Additional Visualisation
(a) Model 1:Cora, Model 2:Pubmed
(b) Overlap with GnnMerge++
(c) Overlap with WAvg.
@@
(a) Model 1:Pubmed, Model 2:WikiCS
(b) Overlap with GnnMerge++
(c) Overlap with WAvg.
@@
(a) Model 1:Citeseer, Model 2:WikiCS
(b) Overlap with GnnMerge++
(c) Overlap with WAvg.
@@
Dataset M Raw WAvg. Permute ZipIt! Surgery GnnMerge GnnMerge++
AmzComp GraphSage 1 95.46 58.55 84.47 66.92 85.88 93.89 94.11
GraphSage 2 92.83 75.91 69.14 74.74 69.92 91.01 91.51
NodeFormer 1 93.33 49.15 - - 87.86 90.22 89.42
NodeFormer 2 91.96 66.51 - - 80.85 88.85 86.71
WikiCS GraphSage 1 86.64 80.36 83.35 76.00 84.14 84.09 84.43
GraphSage 2 84.99 76.80 56.85 64.50 82.23 83.73 83.56
NodeFormer 1 79.21 56.12 - - 76.11 79.02 76.57
NodeFormer 2 78.54 70.19 - - 71.34 75.92 72.43
[…]

### v2

@@ B.4 Baselines
Dataset #Nodes #Edges #Classes #Features
Cora (Yang et al., 2016) 2,708 5,429 7 1,433
Citeseer (Yang et al., 2016) 3,312 4,732 6 3,703
Pubmed (Yang et al., 2016) 19,717 44,338 3 500
Arxiv (Hu et al., 2020) 169,443 2,315,598 40 128
WikiCS (Mernyei & Cangea, 2020) 11,701 431,726 10 300
AmzPhoto (Shchur et al., 2018) 7,650 238,162 8 745
AmzComp (Shchur et al., 2018) 13,752 491,722 10 767
Reddit (Hamilton et al., 2017a) 232,965 114,615,892 41 602
Table 4: Datasets used for benchmarking GnnMerge.
@@ B.5 Additional Experimental Details
Dataset M Raw WAvg. Permute ZipIt! Surgery GnnMerge GnnMerge++
AmzComp GraphSage 1 95.46 58.55 84.47 66.92 85.88 93.89 94.11
GraphSage 2 92.83 75.91 69.14 74.74 69.92 91.01 91.51
NodeFormer 1 93.33 49.15 - - 87.86 90.22 89.42
NodeFormer 2 91.96 66.51 - - 80.85 88.85 86.71
WikiCS GraphSage 1 86.64 80.36 83.35 76.00 84.14 84.09 84.43
GraphSage 2 84.99 76.80 56.85 64.50 82.23 83.73 83.56
NodeFormer 1 79.21 56.12 - - 76.11 79.02 76.57
NodeFormer 2 78.54 70.19 - - 71.34 75.92 72.43
[…]
Arch. Datasets Raw WAvg. Permute ZipIt! Surgery GnnMerge GnnMerge++
GraphSage Arxiv 74.65 55.13 58.45 59.13 68.39 72.85 72.78
WikiCS 78.82 72.84 56.17 56.68 69.22 78.65 78.63
Arxiv 74.65 68.77 60.00 63.60 67.85 74.24 74.29
Pubmed 77.96 72.32 62.21 65.94 75.12 78.01 77.97
NodeFormer Cora 81.09 50.72 - - 77.61 76.59 75.27
Citeseer 81.35 71.71 - - 79.15 77.83 76.08
Pubmed 80.08 57.60 - - 75.74 79.78 79.10
WikiCS 74.17 62.80 - - 67.48 72.42 69.21
[…]
@@ Appendix C Additional Results and Analysis
Tables 91011 and 12 contain the full results for the Different Datasets experiments, with varying number of models being merged.
@@ C.2 Generalization to Gnn Architectures
We further benchmark GnnMerge on GraphSage and NodeFormer in Tables 6 and 5. The results follow the same trend observed with Gcn on both architectures, and thereby establishing the robustness of GnnMerge to accommodate diverse Gnn architectures.
Tasks Raw Random WAvg. Permute ZipIt! Surgery GnnMerge GnnMerge++
Arxiv-NC 73.10 9.31 57.99 61.03 60.72 66.54 73.03 73.01
Pubmed-LP 97.05 90.22 91.32 91.78 91.50 93.67 96.16 96.23
WikiCS-NC 79.32 13.99 75.18 74.79 75.88 77.02 78.79 78.88
Cora-LP 94.34 83.62 77.28 76.90 80.56 88.39 94.12 94.45
WikiCS-NC 79.32 22.87 69.46 71.30 73.35 76.93 78.88 78.91
Pubmed-LP 97.05 91.50 88.70 90.07 89.45 92.71 96.26 96.37
[…]
C.3 Different Tasks
Table 7 presents results for a more challenging scenario: merging Gcns trained on two distinct tasks—node classification and link prediction—across two different datasets. In most cases, baseline methods perform no better than a randomly initialized Gcn on the link prediction task. Preserving node classification accuracy leads to a severe degradation in link prediction AUC for the baselines. In contrast, GnnMerge and GnnMerge++ achieve accuracies comparable to the individual models on their respective tasks.
C.4 Speed Efficiency
Dataset Scratch Train Time GnnMerge GnnMerge++
Arxiv 24.19s 3.75s 1.67s
Reddit 697.88s 102.99s 5.12s
Table 8: Running times for the In-domain dataset task.
C.5 Data Efficiency
C.5.1 Target Node Sampling
@@
Figure 5a and 5b depict the variation in test accuracies as the percentage of nodes utilized for alignment is varied on the arxiv and reddit datasets.
@@
C.5.2 1-Hop Neighbour Condensation
In an Mpnn architecture, each layer computes node embeddings by aggregating messages from a node’s 1-hop neighbors. As a result, at any given layer, a target node’s embedding depends exclusively on its immediate neighbors, if the input is fixed. This property enables the graph to be reduced to only the target nodes and their 1-hop neighbors, significantly decreasing its size. When combined with node sampling, this leads to a substantial reduction in both memory requirements and merging time. Figures 5c and 5d illustrate the variation in memory consumption across different node sampling levels for the Arxiv and Reddit datasets, respectively. The overall sampling procedure leads to 3 benefits:
@@
(c) Graph Size(Edges) of Arxiv as the sampling ratio is varied
(d) Graph Size(Edges) of Reddit as the sampling ratio is varied